OpenStudy (anonymous):

Add the complex expressions, and simplify. (1 – 3.5i) + (3.6 + 6.5i)

OpenStudy (anonymous):

ok distribute it out first and lets see what you get

OpenStudy (anonymous):

(1+3.6) - (3.5i+6.5i)

OpenStudy (anonymous):

nope here is how you distribute \[(a+b) * (c + d) = ac + ad + bc + bd\]

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

yes I am sorry I am helping others and keeping up with my own

OpenStudy (anonymous):

1-3.5i + 3.6 + 6.5i = 1+3.6 *1+6.5i= 3.7 + 6.6i = 3.5i+3.6 * 3.5i+6.5i= 7.1i + 10i = 17i does that look right

OpenStudy (anonymous):

or is it confusing?

OpenStudy (anonymous):

I will tell you the answer is 4.6 + 3i

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lets double check your FOIL

OpenStudy (anonymous):

thank you was it that confusing?

OpenStudy (anonymous):

All you do is add the real parts and then add the imaginary parts.. There shouldnt be any distributing

OpenStudy (anonymous):

You FOIL is very confusing haha.

OpenStudy (anonymous):

works lets double check

OpenStudy (anonymous):

Im really really sorry!!!!! I thought it was (1 – 3.5i) (3.6 + 6.5i) and not (1 – 3.5i) + (3.6 + 6.5i)

OpenStudy (anonymous):

tyler is right!

OpenStudy (anonymous):

\[(1 + 3.6) + (-3.5 + 6.5)i\]\[4.6 + 3i\]

OpenStudy (anonymous):

okay thank you tyler

OpenStudy (anonymous):

and you too Walleye